OpenStudy (anonymous):

help ! below ... Write the series in summation notation: -3 + 6 - 9 + 12 - 15

OpenStudy (anonymous):

a) \[\sum_{k=0}^{4} \left( -1 \right)^{k} -1(3k)\] b) \[\sum_{k=1}^{5} \left( -1 \right)^{k} (3k)\] c) \[\sum_{k=1}^{5} \left( -1 \right)^{k} -1(3k)\] d) \[\sum_{k=0}^{4} \left( -1 \right)^{k} (3k)\]

OpenStudy (fifciol):

b

OpenStudy (anonymous):

thanks @Fifciol